Roll 4 dice. what is the probability of getting at least two ones?

Assume the dices are fair.
Answer #1

Binomial distribution: P(X) = nCx px qn-x

P(getting a 1 on a die roll), p = 1/6

q = 1 - p = 5/6

Number of dice rolled, n = 4

P(getting at least 2 ones) = 1 - P(getting less than 2 ones)

= 1 - P(getting zero 1s) - P(getting one 1)

= 1 - (5/6)4 - (4x(1/6)x(5/6)3)

= 1 - 0.4823 - 0.3858

= 0.1319